# Transformer Multivariate Forecasting: Less is More?

### 30 Dec 2023 v1

@@ Abstract
In the domain of multivariate forecasting, transformer models stand out as powerful apparatus, displaying exceptional capabilities in handling messy datasets from real-world contexts. However, the inherent complexity of these datasets, characterized by numerous variables and lengthy temporal sequences, poses challenges, including increased noise and extended model runtime. This paper focuses on reducing redundant information to elevate forecasting accuracy while optimizing runtime efficiency. We propose a novel transformer forecasting framework enhanced by Principal Component Analysis (PCA) to tackle this challenge. The framework is evaluated by five state-of-the-art (SOTA) models and four diverse real-world datasets. Our experimental results demonstrate the framework’s ability to minimize prediction errors across all models and datasets while significantly reducing runtime. From the model perspective, one of the PCA-enhanced models: PCA+Crossformer, reduces mean square errors (MSE) by 33.3% and decreases runtime by 49.2% on average. From the dataset perspective, the framework delivers 14.3% MSE and 76.6% runtime reduction on Electricity datasets, as well as 4.8% MSE and 86.9% runtime reduction on Traffic datasets. This study aims to advance various SOTA models and enhance transformer-based time series forecasting for intricate data.
@@ Introduction
Informer (Zhou et al. 2021) and Autoformer (Wu et al. 2021) focuses on adapting components for temporal long sequences. The Crossformer model (Zhang and Yan 2022) presents its capability to preserve time and dimension information while effectively capturing cross-time and cross-dimension dependencies, enhancing its performance in multivariate time series forecasting. The Non-stationary Transformer (Liu et al. 2022)), on the other hand, unifies input, converts output, and solves the over-stationary problem to forecast results. PatchTST (Nie et al. 2022) utilizes patching and channel-independent architecture. It facilitates the model to capture local semantic information and longer lookback windows. iTransformer (Liu et al. 2023) inverts the structure of the transformer model without modify components to enhance the performance of forecasting. It provides an alternative architecture for time series forecasting. Additionally, many recent transformer-based models such as Scaleformer (Shabani et al. 2022), TDformer (Zhang et al. 2022), GCformer (Zhao et al. 2023), PDTrans (Tong, Xie, and Zhang 2023) ect, offering diverse approaches to time series forecasting. A comprehensive survey by (Wen et al. 2022) explores transformer models for time series forecasting, which provides some valuable insights. Besides transformer models, researchers also adopt other techniques to address time series forecasting, such as graph representation learning (Jin et al. 2022) and Large Language Models (LLMs) (Jin et al. 2023).
@@
To evaluate all transformer models, we adopt four long-term time series datasets widely accepted as standard data sources. They are ETTh1, Weather, Traffic, and Electricity. The targeted or predicted variables for these datasets are oil temperature, wet bulb, 862th sensor, and 321th client, respectively. Furthermore, these four datasets contain different ranges of variables, varying from 6 to 861 with large time steps. Having dataset with large variables and long time steps can ensure stable results with less over-fitting. Table 2 highlights a summary of information about four datasets. A detailed description of variables for these datasets can be found in (Zhou et al. 2021; Wu et al. 2021).
@@
Table 2: Summary of information about four datasets (targets are not counted into variables).
(MSE normal-$\mid$ Time) (MSE normal-$\mid$ Time) (MSE normal-$\mid$ Time) (MSE normal-$\mid$ Time) (MSE normal-$\mid$ Time) (MSE normal-$\mid$ Time)
Table 3: MSE and runtime reductions of PCA-enhanced transformer forecasting models (PCA+N.S. Trans.: PCA + Non-stationary Transformer). This table is the summary of the accuracy result Table 6 and corresponding elapsed time result Table 7.
@@
All experiments are implemented on a platform with a single Nvidia TU02 GPU. We test all transformer-based models (refer to Table 1), currently considered to be SOTA for time series forecasting. For each experiment, all models are configured with pre-selected default hyperparameters specific to each dataset for their best performance. The prediction length is set to 96 (based on the randomly selection) for all experiments as the long-term forecasting setting. For the models’ evaluation metrics, we use mean squared error (MSE) and mean absolute error (MAE). The code [^1] We will publish the code and scripts after notification of paper acceptance. in this paper is constructed based on the Time Series Library by (Wu et al. 2023).
@@
Dataset Total PCA Information Dataset Best Result Promotion Model
Variables Components Kept (PCA) Ratio ( MSE $\mid$ Time) ( MSE $\mid$ Time)
ETTh1 6 2 70.1% 33.3% 0.05561 71 0.2% 4.1% PCA+Patch.
Weather 20 2 63.1% 10.0% 0.00124 653 8.3% -4.5% PCA+iTrans.
Electricity 320 80 94.7% 25.0% 0.23528 925 14.3% 76.6% PCA+Crossf.
Traffic 861 1 57.6% 0.1% 0.14789 2590 4.8% 86.9% PCA+Crossf.
Table 4: Information Kept Ratio by PCA with the Best Performance Model with Most less MSE on each Dataset.
@@
Table 3 summarises experiment results with PCA-enhanced and non-PCA-enhanced methods across all four datasets. The detailed results for accuracy and running time is in Appendix (Table 6 and Table 7). The proposed PCA-enhanced method exhibits significant reduction MSE and running time for all models except iTransformer. The results highlight that we can improve accuracy and efficiency by reducing input data dimensions. The entire process emphasizes the principle of “less is more”. Based on Table 3, we can summarize the experiment results as follows:
@@
Dataset Total PCA Information Dataset
Variables Components Kept (PCA) Ratio
@@
Figure 4: Left: PCC of ETTh1 dataset. Right: PCC of Weather dataset.
@@ Acknowledgments
Appendix A Appendix
The following contents are all Appendix.
Get Table 3 from Table 6 and Table 7.
Table 3 serves as a consolidation of accuracy results from Table 6 and corresponding runtime results from Table 7. Each entry in Table 3 represents the best-performing (lowest MSE, bold with blue cell color) model from Table 6, along with its associated runtime, in comparison with the original (w/o PCA) model and its corresponding time. For instance, the optimal performance of PCA+PatchTST on the Electricity dataset in Table 6 is 0.30114 (lowest MSE, bold with blue cell color), compared to 0.60771 (w/o PCA). Consequently, the reduction in MSE for PCA+PatchTST on the electricity dataset is 2.14%, as reflected in Table 3.
Table 6: Accuracy result of PCA-enhanced transformer forecasting models (PCA+N.S. Trans.: PCA + Non-stationary Transformer).
Dataset Variables PCA PCA+ PCA+ PCA+ PCA+N.S. PCA+ PCA+
2 71 241 310 118 87 219
4 64 246 311 141 84 126
ETTh1 6 w/o PCA 74 252 310 177 88 126
2 231 1915 686 415 653 375
5 241 1264 690 414 538 377
10 250 1399 701 416 603 379
15 283 1224 689 418 440 378
Weather 20 w/o PCA 311 1768 685 417 625 378
2 221 1430 975 876 200 283
20 246 883 723 1198 209 241
40 352 904 604 1043 207 241
80 579 925 733 887 205 245
160 1038 1620 748 1394 216 250
240 1510 2847 622 1418 380 303
Electricity 320 w/o PCA 1953 3960 759 1768 287 303
1 667 2590 320 146 346 135
2 735 2672 322 149 378 136
25 716 2395 321 152 162 138
50 726 2670 330 153 329 139
105 743 2760 330 155 358 141
215 1285 6122 332 159 409 145
430 2398 10330 348 167 683 154
645 3554 10171 353 177 755 162
Traffic 861 w/o PCA 4626 19749 354 173 625 154
Table 7: Runtime result of PCA-enhanced transformers forecasting model (unit: second).

### 07 Mar 2024 v2

@@ Abstract
In the domain of multivariate forecasting, transformer models stand out as powerful apparatus, displaying exceptional capabilities in handling messy datasets from real-world contexts. However, the inherent complexity of these datasets, characterized by numerous variables and lengthy temporal sequences, poses challenges, including increased noise and extended model runtime. This paper focuses on reducing redundant information to elevate forecasting accuracy while optimizing runtime efficiency. We propose a novel transformer forecasting framework enhanced by Principal Component Analysis (PCA) to tackle this challenge. The framework is evaluated by five state-of-the-art (SOTA) models and four diverse real-world datasets. Our experimental results demonstrate the framework’s ability to minimize prediction errors across all models and datasets while significantly reducing runtime. From the model perspective, one of the PCA-enhanced models: PCA+Crossformer, reduces mean square errors (MSE) by 33.3% and decreases runtime by 49.2% on average. From the dataset perspective, the framework delivers 14.3% MSE and 76.6% runtime reduction on Electricity datasets, as well as 4.8% MSE and 86.9% runtime reduction on Traffic datasets. This study aims to advance various SOTA models and enhance transformer-based time series forecasting for intricate data. Code is available at: https://github.com/jingjing-unilu/PCA˙Transformer
@@ Introduction
Informer (Zhou et al. 2021) and Autoformer (Wu et al. 2021) focuses on adapting components for temporal long sequences. The Crossformer model (Zhang and Yan 2022) presents its capability to preserve time and dimension information while effectively capturing cross-time and cross-dimension dependencies, enhancing its performance in multivariate time series forecasting. The Non-stationary Transformer (Liu et al. 2022)), on the other hand, unifies input, converts output, and solves the over-stationary problem to forecast results. PatchTST (Nie et al. 2022) utilizes patching and channel-independent architecture. It facilitates the model to capture local semantic information and longer lookback windows. iTransformer (Liu et al. 2023) inverts the structure of the transformer model without modify components to enhance the performance of forecasting. It provides an alternative architecture for time series forecasting. Additionally, many recent transformer-based models such as Scaleformer (Shabani et al. 2022), TDformer (Zhang et al. 2022), GCformer (Zhao et al. 2023), PDTrans (Tong, Xie, and Zhang 2023) etc, offering diverse approaches to time series forecasting. In addition, (WU et al. 2024a, b) introduces credit default swap (CDS) dataset for testing transformer models. A comprehensive survey by (Wen et al. 2022) explores transformer models for time series forecasting, which provides some valuable insights. Besides transformer models, researchers also adopt other techniques to address time series forecasting, such as graph representation learning (Jin et al. 2022) and Large Language Models (LLMs) (Jin et al. 2023).
@@
To evaluate all transformer models, we adopt four long-term time series datasets widely accepted as standard data sources. They are ETTh1, Weather, Traffic, and Electricity. The targeted or predicted variables for these datasets are oil temperature, wet bulb, 862th sensor, and 321th client, respectively. Furthermore, these four datasets contain different ranges of variables, varying from 7 to 862 with large time steps. Having dataset with large variables and long time steps can ensure stable results with less over-fitting. Table 2 highlights a summary of information about four datasets. A detailed description of variables for these datasets can be found in (Zhou et al. 2021; Wu et al. 2021).
@@
Table 2: Summary of information about four datasets (targets are counted into variables).
@@
Table 3: Accuracy results of transformer forecasting models (w/o PCA) and PCA-enhanced transformer forecasting models. Note: 1) w/o PCA: without PCA; 2) PCA+N.S. Trans.: PCA + Non-stationary Transformer.
Dataset Variables PCA PCA+ PCA+ PCA+ PCA+N.S. PCA+ PCA+
2 71 241 310 118 87 219
4 64 246 311 141 84 126
ETTh1 6 w/o PCA 74 252 310 177 88 126
2 231 1915 686 415 653 375
5 241 1264 690 414 538 377
10 250 1399 701 416 603 379
15 283 1224 689 418 440 378
Weather 20 w/o PCA 311 1768 685 417 625 378
2 221 1430 975 876 200 283
20 246 883 723 1198 209 241
40 352 904 604 1043 207 241
80 579 925 733 887 205 245
160 1038 1620 748 1394 216 250
240 1510 2847 622 1418 380 303
Electricity 320 w/o PCA 1953 3960 759 1768 287 303
1 667 2590 320 146 346 135
2 735 2672 322 149 378 136
25 716 2395 321 152 162 138
50 726 2670 330 153 329 139
105 743 2760 330 155 358 141
215 1285 6122 332 159 409 145
430 2398 10330 348 167 683 154
645 3554 10171 353 177 755 162
Traffic 861 w/o PCA 4626 19749 354 173 625 154
Table 4: Runtime results of transformer forecasting models (w/o PCA) and PCA-enhanced transformers forecasting model (unit: second).
@@
All experiments are implemented on a platform with a single Nvidia TU02 GPU. We test all transformer-based models (refer to Table 1), currently considered to be SOTA for time series forecasting. For each experiment, all models are configured with pre-selected default hyperparameters specific to each dataset for their best performance. The prediction length is set to 96 (based on the randomly selection) for all experiments as the long-term forecasting setting. For the models’ evaluation metrics, we use mean squared error (MSE) and mean absolute error (MAE). The code [^1] https://github.com/jingjing-unilu/PCA˙Transformer in this paper is constructed based on the Time Series Library by (Wu et al. 2023).
@@
w/o Target (MSE normal-$\mid$ Time) (MSE normal-$\mid$ Time) (MSE normal-$\mid$ Time) (MSE normal-$\mid$ Time) (MSE normal-$\mid$ Time) (MSE normal-$\mid$ Time)
Table 5: MSE and runtime reductions of PCA-enhanced transformer forecasting models. This table is the summary of the accuracy result Table 3 and corresponding runtime result Table 4. The baselines are corresponding original transformer models.
@@
Table 5 summarises experiment results with PCA-enhanced and non-PCA-enhanced methods across all four datasets. Each entry in Table 5 represents the best-performing model from Table 3 and its runtime Table 4 (lowest MSE with its runtime, bold with blue cell color). Table 5 shows that the proposed PCA-enhanced method exhibits significant reduction MSE and running time for all models except iTransformer. The results highlight that we can improve accuracy and efficiency by reducing input data dimensions. The entire process emphasizes the principle of “less is more”. Based on Table 5, we can summarize the experiment results as follows:
@@
Figure 4: Left: PCC of ETTh1 dataset. Right: PCC of Weather dataset.
Dataset Variables PCA Information Dataset Best Result Best Reduction Model
w/o Target Components Kept (PCA) Ratio ( MSE $\mid$ Time) ( MSE $\mid$ Time)
ETTh1 6 2 70.1% 33.3% 0.05561 71 0.2% 4.1% PCA+Patch.
Weather 20 2 63.1% 10.0% 0.00124 653 8.3% -4.5% PCA+iTrans.
Electricity 320 80 94.7% 25.0% 0.23528 925 14.3% 76.6% PCA+Crossf.
Traffic 861 1 57.6% 0.1% 0.14789 2590 4.8% 86.9% PCA+Crossf.
Table 6: Information Kept Ratio by PCA with the Best Performance Model (Lowest MSE on each Dataset). The baselines are corresponding original transformer models.
Dataset Variables PCA Information Dataset
w/o Target Components Kept (PCA) Ratio